# Cross-Modal Few-Shot Learning: a Generative Transfer Learning Framework

### v1

@@ 5 Conclusions
In this work, we tackled the limitations of unimodal few-shot learning by introducing the cross-modal few-shot learning task, which addresses real-world scenarios involving multiple visual modalities with only a few labeled examples. Unlike classical few-shot learning, our task presents additional challenges due to the inherent variability in visual characteristics, structural properties, and domain gaps between modalities.
To overcome these challenges, we proposed the generative transfer learning framework, designed to enable efficient knowledge transfer from abundant unimodal data to data-scarce multi-modal scenarios. By estimating shared latent concepts from unimodal data and generalizing them to unseen modalities, the GTL framework effectively disentangles modality-independent representations from in-modality disturbances. Our experimental results demonstrated that our GTL framework significantly outperforms state-of-the-art methods on four multi-modal datasets: Sketchy, TU-Berlin, Mask1K, and SKSF-A.
Limitations:
While our approach demonstrates strong performance, the lack of diverse visual multi-modal datasets for visual recognition remains a significant challenge. Most existing datasets focus on a limited number of visual modalities (e.g., RGB and sketch), restricting the evaluation of models designed to handle more complex visual data. Future work will require more extensive and varied multi-modal datasets to fully explore and validate the potential of cross-modal few-shot learning in a broader range of visual contexts.
@@ A.1 Relation to Unimodal Few-shot Learning
In general, unimodal learning tasks can be broadly categorized into three types based on the availability and quantity of labeled data: (i) Supervised Learning, where a large amount of labeled data is available for training, allowing models to learn features and perform accurate recognition (Yan et al., 2023a; Han et al., 2023; Li et al., 2023a); (ii) Few-shot Learning, where only a limited number of labeled samples are provided for each class, challenging the model to generalize effectively from minimal data (Chen et al., 2019; Luo et al., 2023); and (iii) Zero-shot Learning, where no labeled examples are available for certain classes (Wei et al., 2023; Li et al., 2024a; Mirza et al., 2024).
@@
Specifically, Few-shot learning (FSL) encompasses a training phase where a model is trained on a relatively large dataset and an adaptation phase in which the trained model is adjusted to previously unseen tasks with limited labeled samples. Most existing FSL tasks utilize unimodal datasets for training and testing, including popular benchmarks such as ImageNet (Deng et al., 2009), CIFAR (Oreshkin et al., 2018), CUB-200-2021 (Wah et al., 2011), and Stanford Dogs (Khosla et al., 2011). FSL typically involves three main approaches: (i) Meta learning (Sun et al., 2019; Ma et al., 2024), or learning to learn, which optimizes model parameters across diverse learning tasks to enable rapid adaptation to new challenges; (ii) Data-centric learning (Li et al., 2020; Meng et al., 2023; Ma et al., 2024), which focuses on metric learning to compare distances between samples or expand synthetic data facing with data-scarce scenarios; (iii) Transfer learning (Tian et al., 2020; Luo et al., 2023; Zhang et al., 2024) where models pre-trained on large-scale datasets are fine-tuned on few-shot tasks to improve performance to improve performance by leveraging learned representations for more efficient adaptation.
@@
The adaption-based methods (Sun et al., 2019; Ma et al., 2024) are densely connected to the model design, which attempts to directly establish a mapping function between input and prediction. By rapidly updating parameters on new tasks with a small number of samples, these methods facilitate the transfer of knowledge from previously learned tasks, making them highly effective in few-shot learning scenarios.
@@
The data-centric methods utilize synthetic data (Meng et al., 2023) or metric learning (Ma et al., 2024) to adapt data-insufficient scenarios. The former involves generation methods like Generative Adversarial Networks (GAN) (Li et al., 2020) and auto-encoders (Yan et al., 2023b). These approaches generate additional training data to enhance the model’s performance, mitigating the scarcity of labeled examples and improving generalization to new tasks. The latter, metric learning, tries to build data connections with the thoughts of nearby neighbors, focusing on learning a distance metric that clusters similar examples together and separates dissimilar ones. These data-centric methods emphasize the properties of the data to enhance the model’s ability to generalize from a few examples.
@@
The fine-tuning-based methods (Tian et al., 2020; Luo et al., 2023; Zhang et al., 2024) involve using pre-trained models on large datasets and fine-tuning them on the target task with a small number of examples. This approach leverages the knowledge gained from the pre-training phase and transfers it to the specific requirements of the new task. However, with the increasing amount of raw data on the Internet, it’s challenging for pre-trained models, including Vision-Language Pretraining Models (VLM) (Zhu et al., 2024), to generalize to specific novel data, especially when the data is in different modalities.
@@
Recent works on cross-domain learning (Wang et al., 2020; Li et al., 2022; Xu et al., 2023) also focus exclusively on learning from unimodal data. However, their limitations become apparent in complex real-world applications that often require understanding multiple modalities simultaneously. In this work, we extend its applicability of few-shot learning to real-world scenarios where data come from diverse visual modalities.
@@ A.2 Relation to Multi-modal Few-shot Learning
Existing multi-modal few-shot learning methods aim to leverage information from multiple modalities, such as combining visual data with textual (Xing et al., 2019; Tsimpoukelli et al., 2021; Alayrac et al., 2022; Lin et al., 2023c; Shao et al., 2024), audio (Meshry et al., 2021; Majumder et al., 2022; Kong et al., 2024), or tabular data Liu et al. (2022); Ye et al. (2024); Han et al. (2024). These approaches focus on enriching the feature space by integrating heterogeneous data sources, thereby improving the generalization capabilities of few-shot models in scenarios where additional modality information is available. However, most of these methods concentrate on leveraging modalities that are inherently different from visual data, like text or audio, to provide semantic context or supplementary cues. They often assume access to richly annotated data in these auxiliary modalities, which may not always be feasible in practical applications. These methods enhance recognition by supplementing visual features with external information but do not address the challenges that arise when dealing with multiple visual modalities.
Few methods address the challenges of visual multi-modal scenarios where multiple visual modalities of the same object exist, such as RGB images, sketches, infrared images, or depth maps. Visual modalities often exhibit unique characteristics and structural properties, leading to significant domain gaps that complicate direct knowledge transfer between them. Some approaches Bhunia et al. (2022) utilize additional visual modalities to enhance recognition performance on a primary modality; however, they typically treat the additional visual modality as auxiliary information to support unimodal recognition rather than fully integrating multiple visual modalities into a unified learning framework.

### v2

@@ 5 Conclusions
This paper introduces the cross-modal few-shot learning task, which addresses real-world scenarios involving multiple visual modalities with only a few labeled examples. Unlike classical few-shot learning, our task presents additional challenges due to the inherent variability in visual characteristics, structural properties, and domain gaps between modalities. To overcome these challenges, we propose the generative transfer learning framework, which enables efficient knowledge transfer from abundant unimodal data to data-scarce multi-modal scenarios by estimating shared latent concepts from unimodal data and generalizing them to unseen modalities. Extensive experiments on seven unique multi-modal datasets, demonstrated the superiority of the proposed framework by effectively disentangling modality-independent representations from in-modality disturbances. The results also suggest that the model can estimate latent concepts from vast unimodal data and generalize these concepts to unseen modalities using only a few available samples, much like human cognitive processes.
@@ A.1 Relation to Unimodal Few-shot Learning
In general, unimodal learning tasks can be broadly categorized into three types based on the availability and quantity of labeled data: (i) Supervised Learning, where a large amount of labeled data is available for training, allowing models to learn features and perform accurate recognition [69, 16, 30]; (ii) Few-shot Learning, where only a limited number of labeled samples are provided for each class, challenging the model to generalize effectively from minimal data [5, 40]; and (iii) Zero-shot Learning, where no labeled examples are available for certain classes [64, 31, 45].
@@
Specifically, Few-shot learning (FSL) encompasses a training phase where a model is trained on a relatively large dataset and an adaptation phase in which the trained model is adjusted to previously unseen tasks with limited labeled samples. Most existing FSL tasks utilize unimodal datasets for training and testing, including popular benchmarks such as ImageNet [8], CIFAR [48], CUB-200-2021 [61], and Stanford Dogs [24]. FSL typically involves three main approaches: (i) Meta learning [55, 41], or learning to learn, which optimizes model parameters across diverse learning tasks to enable rapid adaptation to new challenges; (ii) Data-centric learning [29, 43, 41], which focuses on metric learning to compare distances between samples or expand synthetic data facing with data-scarce scenarios; (iii) Transfer learning [57, 40, 75] where models pre-trained on large-scale datasets are fine-tuned on few-shot tasks to improve performance to improve performance by leveraging learned representations for more efficient adaptation.
@@
The meta-learning methods [55, 4, 41, 12] are densely connected to the model design, which attempts to directly establish a mapping function between input and prediction. Cao et al. [4] separate targets as fine-grained parts to build concept learners. Ma et al. [41] incorporate semantic information and spatial position into a reconstruction framework to strengthen the feature alignment between seen and unseen samples. By rapidly updating parameters on new tasks with a small number of samples, these methods facilitate the transfer of knowledge from previously learned tasks, making them highly effective in few-shot learning scenarios.
@@
The data-centric methods utilize synthetic data [43] or metric learning [41] to adapt data-insufficient scenarios. The former involves generation methods like Generative Adversarial Networks (GAN) [29] and auto-encoders [70]. These approaches generate additional training data to enhance the model’s performance, mitigating the scarcity of labeled examples and improving generalization to new tasks. The latter, metric learning, tries to build data connections with the thoughts of nearby neighbors, focusing on learning a distance metric that clusters similar examples together and separates dissimilar ones. These data-centric methods emphasize the properties of the data to enhance the model’s ability to generalize from a few examples.
@@
The fine-tuning-based methods [57, 40, 75] involve using pre-trained models on large datasets and fine-tuning them on the target task with a small number of examples. This approach leverages the knowledge gained from the pre-training phase and transfers it to the specific requirements of the new task. However, with the increasing amount of raw data on the Internet, it’s challenging for pre-trained models, including Vision-Language Pretraining Models (VLM) [78], to generalize to specific novel data, especially when the data is in different modalities.
@@
Recent works on cross-domain learning [62, 33, 68] also focus exclusively on learning from unimodal data. However, their limitations become apparent in complex real-world applications that often require understanding multiple modalities simultaneously. In this work, we extend its applicability of few-shot learning to real-world scenarios where data come from diverse visual modalities.
@@ A.2 Relation to Multi-modal Few-shot Learning
Existing multi-modal few-shot learning methods aim to leverage information from multiple modalities, such as combining visual data with textual [67, 58, 1, 37, 52], audio [44, 42, 26], or tabular data [39, 71, 15]. These approaches focus on enriching the feature space by integrating heterogeneous data sources, thereby improving the generalization capabilities of few-shot models in scenarios where additional modality information is available. However, most of these methods concentrate on leveraging modalities that are inherently different from visual data, like text or audio, to provide semantic context or supplementary cues. They often assume access to richly annotated data in these auxiliary modalities, which may not always be feasible in practical applications. These methods enhance recognition by supplementing visual features with external information but do not address the challenges that arise when dealing with multiple visual modalities.
Few methods address the challenges of visual multi-modal scenarios where multiple visual modalities of the same object exist, such as RGB images, sketches, infrared images, or depth maps. Visual modalities often exhibit unique characteristics and structural properties, leading to significant domain gaps that complicate direct knowledge transfer between them. Some approaches Bhunia et al. [3] utilize additional visual modalities to enhance recognition performance on a primary modality; however, they typically treat the additional visual modality as auxiliary information to support unimodal recognition rather than fully integrating multiple visual modalities into a unified learning framework.